# More is uncorrelated: Tuning the local correlations of SU($N$) Fermi-Hubbard systems via controlled symmetry breaking

### v1

@@ IV.3 Broken SU(4) symmetry
Figure 7: Two-flavors entropy $s_{\alpha\beta}$, single-flavor entropy $s_{\alpha},s_{\beta}$ and inter-flavor mutual information $I_{\alpha\beta}$, plotted against the interaction strength $U$, for different values of the Raman coupling $\Omega$, indicated on top of each column, for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel). The insets in the bottom panel are close-ups of the mutual information between Raman-coupled flavors, highlighting that it vanishes for $U>U^{\mathrm{R}}_{\mathrm{c}}(\Omega)$. Raman-free flavors display a clear crossover between the SU($2$) and the SU($4$) ($\Omega=0$) model as $\Omega$ increases. All quantities are in units of $\log(2)$.
We now include a finite $\Omega\neq 0$. As a result the original SU($4$) symmetry, supplemented by a global charge conservation is reduced to $SU(2)\otimes U(1)\otimes U(1)$, where SU($2$) is the usual symmetry for the Raman-free flavors, and the two U($1$) are generated by the conservation of $\tilde{N}_{3}-\tilde{N}_{4}$ and $N_{1}+N_{2}-\tilde{N}_{3}-\tilde{N}_{4}$ respectively.
@@
The evolution of the average double occupancies per site across the Mott transition, displayed in Fig. 6, reflects the above-mentioned polarization of the Raman-coupled flavors, as well as the SU(44)-SU(22) crossover for the Raman-free ones. For the latter case, shown in the top panel of Fig. 6, when $\Omega=0$ the double occupancies decrease and approach $1/6$ in the limit $U\gg D$, as expected for a SU($4$) Mott insulator. In contrast, for any $\Omega>0$, double occupancies are suppressed in the Mott insulating phase, aligning with the SU($2$) behavior as the coupling with the Raman field increases. This clearly highlights the presence of a crossover between the two models driven by the tuning of $\Omega$. On the other hand, the double occupancies of the Raman-coupled flavors, shown in the bottom panel of Fig. 6, display clear signs of growing polarization between the two flavors as $U$ increases. When $U$ increases beyond $U_{p}\left(\Omega\right)$, the double occupancies drop to zero, signaling that the Raman-coupled flavors have become fully polarized, with only one of their two bands occupied.
@@
As shown in the top panel of Fig. 6, the quasiparticle weight of the Raman-free flavors vanishes at $U_{c}\left(\Omega\right)$, marking the transition into the Mott insulating phase. For every $\Omega$, $Z$ lies between its values observed in the SU($4$) and in the SU($2$) model, approaching the SU($2$) curve as $\Omega/D$ grows closer to $1$. Indeed, for these flavors the critical interaction strength for the metal-to-insulator transition decreases from $U_{\mathrm{c}}\simeq 4.2$, when $\Omega=0$ (characteristic of the SU($4$) model) to $U_{\mathrm{c}},\simeq 2.8$, when $\Omega\simeq D$, which is precisely the critical value observed in the SU($2$) model. On the contrary, the quasiparticle weight of the Raman-coupled flavors tends to 1 for large values of $U/D$, for all $\Omega>0$, as shown in the bottom panel of Fig. 6. This reflects the hybrid nature of the large-$U$ state, which is not a fully incoherent Mott insulator but has the Raman-coupled flavors behaving as in a band insulator, where quasiparticles remain well-defined [^1] For the Raman-coupled flavors, $Z$ was not computed using Eq. 11, but rather as the residue of the pole of the Green’s function at $\omega^{*}$, determined by the condition $\omega^{*}-\varepsilon-\text{Re}\Sigma(\omega^{*})=0$. Due to the locality of the self-energy, $Z$ is independent on $\varepsilon$ , therefore we compute it for $\varepsilon=0$. This approach is made necessary because, in a band insulator, quasiparticles are not located at the Fermi level but in the conduction and valence bands..
@@
Importantly, we can study how inter-flavor correlations evolve with the symmetry breaking term, as shown in Fig. 7. At $\Omega=0$, when $U>U_{\mathrm{c}}\left(0\right)$ the system is a weakly correlated SU($4$) Mott insulator, as already discussed in Sec. IV.1 (see Fig. 3). As $\Omega$ increases, correlations between Raman-free flavors in the insulating phase become significantly stronger, eventually matching the values observed in the SU($2$) model, as shown in the top row of Fig. 7. It is then clear how reducing the symmetry between flavors leads to an enhancement of local inter-particle correlations, recovering strong SU(2)-like correlations in the Mott-insulating state.
@@
The evolution of entropy-based quantities not only neatly captures the crossover between the SU($4$) and the SU($2$) Mott-Hubbard transitions, but also illustrates how inter-particle correlations in the Mott phase can be effectively tuned by controlling the symmetry of the model. Remarkably, one can recover a strongly-correlated insulator from a multi-component system by breaking the underlying flavor symmetry.
@@ IV.4 $\Omega-U$ phase diagram and the tricritical point
The transition between these phases is determined by the two critical values of $U$. The red line in Fig. 8 represents $U_{\mathrm{c}}\left(\Omega\right)$ and marks the Mott transition of the Raman-free flavors, signaled by the vanishing of the quasiparticle weights $Z_{1}$, $Z_{2}$. The blue line represents $U_{\mathrm{p}}\left(\Omega\right)$ and determines the critical points where the Raman-coupled flavors become fully polarized, meaning that $\tilde{n}_{4}-\tilde{n}_{3}=1$, and form a band insulator. These two transition lines appear to intersect at a tricritical point, $\left(U^{\mathrm{T}},\Omega^{\mathrm{T}}\right)$, where all three phases coexist. At this point, the nature of the transition for the Raman-coupled flavors changes. For $\Omega<\Omega^{\mathrm{T}}$, the transition is found to be clearly first order, characterized by a discontinuous jump in the flavor polarization $\tilde{n}_{4}-\tilde{n}_{3}$, while for $\Omega>\Omega^{\mathrm{T}}$, the transition appears of second order within our numerical accuracy, with the polarization smoothly approaching $1$, i.e. the value defining a complete polarization. This difference is illustrated in the bottom panel of Fig. 8, where both horizontal and vertical cuts through the phase diagram show the polarization behavior. Additionally, for $\Omega<\Omega^{\mathrm{T}}$, the two transition lines merge, indicating that both Raman-free and Raman-coupled flavors undergo the phase transition simultaneously.
@@ V Summary and outlook
In this work, we studied the degree of correlation between flavors in the SU($N$) Hubbard model in the geometrical framework associated to quantum relative entropies, an ideal suit for open quantum subsystems [55, 34, 35, 37, 29, 30], such as fermionic orbitals embedded in a lattice. By recognizing that the symmetries of the model constrain all local (i.e. residing on a single-site) reduced density matrices to be pseudo-classical [38], we ensured that neither entanglement nor quantum discord is shared by the local flavors, so that the inter-flavor mutual information gives a faithful measure of local classical correlations between particles [56], relating to the concept of fermionic nongaussianity [45, 46, 47, 49, 50, 51, 52, 53, 54, 48]. We characterized the behavior of this quantity across the Mott-Hubbard transition for $N=2$ and $N=4$ and examined the effects of a tunable symmetry breaking term in the latter case, effectively uncovering a rich crossover between SU(4) and SU(2)-symmetric case. We benchmarked the behavior of the inter-flavor mutual information, a quantity accessible to cold atom experiments, by comparing it to the evolution of the quasiparticle weight, an experimentally challenging quantity, naturally linked to Fermi liquid theory and its breakdown at genuine Mott-Hubbard transitions.
@@
In the absence of symmetry breaking, we showed how classical local correlations monotonically increase with the inter-flavor interaction strength $U$, eventually saturating in the Mott-Hubbard insulating phase. However, the number of available flavors in the lattice significantly affects the degree of correlation of the insulator, which becomes less correlated as $N$ increases. We formalized this statement in the atomic limit, where the hopping is set to zero and the model, reduced to a collection of Mott-Hubbard states, is analytically solvable. In this regime, we computed the mutual information between flavor pairs, and found that the correlations vanish in the large-$N$ limit. It follows that in the presence of a large number of flavors, the “correlation-driven” Mott insulator becomes effectively uncorrelated. Additionally, we showed that the vanishing of the inter-flavor mutual information can be characterized analytically at half-filling, as $\mathcal{O}(1/2N^{2})$ for $N\rightarrow\infty$. Since the mutual information is an upper bound for all correlation functions between pairs of local operators, its vanishing implies that all local expectation values factorize in the large-$N$ limit. This is consistent with the evolution of local double occupancies, which approach their factorized values when $N\gg 1$, i.e. $\langle n_{\alpha}n_{\beta}\rangle\rightarrow\langle n_{\alpha}\rangle\langle n_{\beta}\rangle$, confirming the emergence of mean-field behavior. This apparently counter-intuitive decoupling is actually a well-explored feature of Yang-Mills theories, namely gauge theories based on the group SU($N$), where it was shown that in the large-$N$ limit quantum fluctuations are damped and mean-field-like correlations emerge [84, 85]. Yet, we highlight the striking quantitative distance between the SU(2) and the SU(4) scenarios, with the latter effectively closer to the large-$N$ limit than to the highly singular SU(2) model.
@@
To further investigate the behavior of correlations in the small-$N$ regime, we showed how introducing a suitable Raman term in the $N=4$ case induces a crossover between the SU($4$) and the SU($2$)-symmetric Hubbard model for the Raman-free flavors, while the Raman-coupled ones become completely polarized. This provides a remarkable example of flavor-selective behavior: the Raman-free flavors undergo an SU($2$) Mott transition, whereas the Raman-coupled ones form an uncorrelated band insulator. The inter-flavor mutual information neatly captures this crossover, and provides clear-cut insight into the contrasting nature of Raman-coupled and Raman-free flavors. Thus, we reinforce previous evidence that this quantity, as well as closely related geometrical measures of entanglement and quantum correlations, represent a crucial tool not only for the generic detection and characterization of critical behavior and phase transitions, but also for the classification of fermionic states in terms of strong, interaction-induced, correlations [56, 29, 30].
We conclude by remarking one more time how this model is readily realizable in experiments with ultracold atoms in optical lattices, which have repeatedly proven their role as essential platforms to probe the physical properties of SU($N$) fermions [11, 12, 13, 15, 10, 14]. As the symmetry-breaking term can be implemented via Raman transitions, our results provide a practical route to directly tune between SU($4$) and SU($2$) Hubbard models, in this class of experiments. Remarkably, the resulting crossover can be characterized thoroughly by means of experimentally accessible observables, such as flavor-resolved densities and double occupancies, providing direct insights into the nature of local correlations, as measured by the inter-flavor mutual information. Furthermore, the possibility to smoothly connect SU($N$)-symmetric models with different $N$ via controlled symmetry-breaking was already hinted-at in [11], highlighting the strong connection between our theoretical results and their realization in state-of-the-art experiments.
@@
Experimental realizations in cold-atom setups offer an intriguing platform to further explore the crossover between the SU(4) and SU(2) models and to potentially realize a quantum tricritical point in practice. To go beyond our present results, it is essential to extend the analysis to include an account of nonlocal correlations in finite dimensions [29, 86] and multipartite entanglement, as witnessed by quench dynamics [24] or measured by the mutual information between many natural orbitals [74]. While the former two avenues will surely strengthen the connection with experimental observations, the latter has the potential to deepen the understanding of the intricate connection between traditional quantum resources and the concept of fermionic nongaussianity [45, 46, 47, 48, 49, 50, 51, 52, 53, 54].

### v2

@@ IV.3 Broken SU(4) symmetry
Figure 7: Two-flavors entropy $s_{\alpha\beta}$, single-flavor entropy $s_{\alpha},s_{\beta}$ and inter-flavor mutual information $I_{\alpha\beta}$, plotted against the interaction strength $U$, for different values of the Raman coupling $\Omega$, indicated on top of each column, for the Raman-free flavors (top panel) and the Raman-coupled flavors (bottom panel). The insets in the bottom panel are close-ups of the mutual information between Raman-coupled flavors, highlighting that it vanishes for $U>U^{\mathrm{R}}_{\mathrm{c}}(\Omega)$. Raman-free flavors display a clear crossover between the SU(22) and the SU(44) ($\Omega=0$) model as $\Omega$ increases. All quantities are in units of $\log(2)$.
We now include a finite $\Omega\neq 0$. As a result the original SU(44) symmetry, supplemented by a global charge conservation is reduced to $SU(2)\otimes U(1)\otimes U(1)$, where SU(22) is the usual symmetry for the Raman-free flavors, and the two U(11) are generated by the conservation of $\tilde{N}_{3}-\tilde{N}_{4}$ and $N_{1}+N_{2}-\tilde{N}_{3}-\tilde{N}_{4}$ respectively.
@@
The evolution of the average double occupancies per site across the Mott transition, displayed in Fig. 6, reflects the above-mentioned polarization of the Raman-coupled flavors, as well as the SU(44)-SU(22) crossover for the Raman-free ones. For the latter case, shown in the top panel of Fig. 6, when $\Omega=0$ the double occupancies decrease and approach $1/6$ in the limit $U\gg D$, as expected for a SU(44) Mott insulator. In contrast, for any $\Omega>0$, double occupancies are suppressed in the Mott insulating phase, aligning with the SU(22) behavior as the coupling with the Raman field increases. This clearly highlights the presence of a crossover between the two models driven by the tuning of $\Omega$. On the other hand, the double occupancies of the Raman-coupled flavors, shown in the bottom panel of Fig. 6, display clear signs of growing polarization between the two flavors as $U$ increases. When $U$ increases beyond $U_{p}\left(\Omega\right)$, the double occupancies drop to zero, signaling that the Raman-coupled flavors have become fully polarized, with only one of their two bands occupied.
@@
As shown in the top panel of Fig. 6, the quasiparticle weight of the Raman-free flavors vanishes at $U_{c}\left(\Omega\right)$, marking the transition into the Mott insulating phase. For every $\Omega$, $Z$ lies between its values observed in the SU(44) and in the SU(22) model, approaching the SU(22) curve as $\Omega/D$ grows closer to 11. Indeed, for these flavors the critical interaction strength for the metal-to-insulator transition decreases from $U_{\mathrm{c}}\simeq 4.2$, when $\Omega=0$ (characteristic of the SU(44) model) to $U_{\mathrm{c}},\simeq 2.8$, when $\Omega\simeq D$, which is precisely the critical value observed in the SU(22) model. On the contrary, the quasiparticle weight of the Raman-coupled flavors tends to 1 for large values of $U/D$, for all $\Omega>0$, as shown in the bottom panel of Fig. 6. This reflects the hybrid nature of the large-$U$ state, which is not a fully incoherent Mott insulator but has the Raman-coupled flavors behaving as in a band insulator, where quasiparticles remain well-defined [^1] For the Raman-coupled flavors, $Z$ was not computed using Eq. 11, but rather as the residue of the pole of the Green’s function at $\omega^{*}$, determined by the condition $\omega^{*}-\varepsilon-\text{Re}\Sigma(\omega^{*})=0$. Due to the locality of the self-energy, $Z$ is independent on $\varepsilon$ , therefore we compute it for $\varepsilon=0$. This approach is made necessary because, in a band insulator, quasiparticles are not located at the Fermi level but in the conduction and valence bands..
@@
Importantly, we can study how inter-flavor correlations evolve with the symmetry breaking term, as shown in Fig. 7. At $\Omega=0$, when $U>U_{\mathrm{c}}\left(0\right)$ the system is a weakly correlated SU(44) Mott insulator, as already discussed in Sec. IV.1 (see Fig. 3). As $\Omega$ increases, correlations between Raman-free flavors in the insulating phase become significantly stronger, eventually matching the values observed in the SU(22) model, as shown in the top row of Fig. 7. It is then clear how reducing the symmetry between flavors leads to an enhancement of local inter-particle correlations, recovering strong SU(2)-like correlations in the Mott-insulating state.
@@
The evolution of entropy-based quantities not only neatly captures the crossover between the SU(44) and the SU(22) Mott-Hubbard transitions, but also illustrates how inter-particle correlations in the Mott phase can be effectively tuned by controlling the symmetry of the model. Remarkably, one can recover a strongly-correlated insulator from a multi-component system by breaking the underlying flavor symmetry.
@@ IV.4 $\Omega-U$ phase diagram and the tricritical point
The transition between these phases is determined by the two critical values of $U$. The red line in Fig. 8 represents $U_{\mathrm{c}}\left(\Omega\right)$ and marks the Mott transition of the Raman-free flavors, signaled by the vanishing of the quasiparticle weights $Z_{1}$, $Z_{2}$. The blue line represents $U_{\mathrm{p}}\left(\Omega\right)$ and determines the critical points where the Raman-coupled flavors become fully polarized, meaning that $\tilde{n}_{4}-\tilde{n}_{3}=1$, and form a band insulator. These two transition lines appear to intersect at a tricritical point, $\left(U^{\mathrm{T}},\Omega^{\mathrm{T}}\right)$, where all three phases coexist. At this point, the nature of the transition for the Raman-coupled flavors changes. For $\Omega<\Omega^{\mathrm{T}}$, the transition is found to be clearly first order, characterized by a discontinuous jump in the flavor polarization $\tilde{n}_{4}-\tilde{n}_{3}$, while for $\Omega>\Omega^{\mathrm{T}}$, the transition appears of second order within our numerical accuracy, with the polarization smoothly approaching 11, i.e. the value defining a complete polarization. This difference is illustrated in the bottom panel of Fig. 8, where both horizontal and vertical cuts through the phase diagram show the polarization behavior. Additionally, for $\Omega<\Omega^{\mathrm{T}}$, the two transition lines merge, indicating that both Raman-free and Raman-coupled flavors undergo the phase transition simultaneously.
@@ V Summary and outlook for experiments
Summary and Conclusions In this work, we studied the degree of correlation between flavors in the SU(NN) Hubbard model in the geometrical framework associated to quantum relative entropies, an ideal suit for open quantum subsystems [55, 34, 35, 37, 29, 30], such as fermionic orbitals embedded in a lattice. By recognizing that the symmetries of the model constrain all local (i.e. residing on a single-site) reduced density matrices to be pseudo-classical [38], we ensured that neither entanglement nor quantum discord is shared by the local flavors, so that the inter-flavor mutual information gives a faithful measure of local classical correlations between particles [56], relating to the concept of fermionic nongaussianity [45, 46, 47, 49, 50, 51, 52, 53, 54, 48]. We characterized the behavior of this quantity across the Mott-Hubbard transition for $N=2$ and $N=4$ and examined the effects of a tunable symmetry breaking term in the latter case, effectively uncovering a rich crossover between SU(4) and SU(2)-symmetric case. We benchmarked the behavior of the inter-flavor mutual information, a quantity accessible to cold atom experiments, by comparing it to the evolution of the quasiparticle weight, an experimentally challenging quantity, naturally linked to Fermi liquid theory and its breakdown at genuine Mott-Hubbard transitions.
@@
In the absence of symmetry breaking, we showed how classical local correlations monotonically increase with the inter-flavor interaction strength $U$, eventually saturating in the Mott-Hubbard insulating phase. However, the number of available flavors in the lattice significantly affects the degree of correlation of the insulator, which becomes less correlated as NN increases. We formalized this statement in the atomic limit, where the hopping is set to zero and the model, reduced to a collection of Mott-Hubbard states, is analytically solvable. In this regime, we computed the mutual information between flavor pairs, and found that the correlations vanish in the large-NN limit. It follows that in the presence of a large number of flavors, the “correlation-driven” Mott insulator becomes effectively uncorrelated. Additionally, we showed that the vanishing of the inter-flavor mutual information can be characterized analytically at half-filling, as $\mathcal{O}(1/2N^{2})$ for $N\rightarrow\infty$. Since the mutual information is an upper bound for all correlation functions between pairs of local operators, its vanishing implies that all local expectation values factorize in the large-NN limit. This is consistent with the evolution of local double occupancies, which approach their factorized values when $N\gg 1$, i.e. $\langle n_{\alpha}n_{\beta}\rangle\rightarrow\langle n_{\alpha}\rangle\langle n_{\beta}\rangle$, confirming the emergence of mean-field behavior. This apparently counter-intuitive decoupling is actually a well-explored feature of Yang-Mills theories, namely gauge theories based on the group SU(NN), where it was shown that in the large-NN limit quantum fluctuations are damped and mean-field-like correlations emerge [84, 85]. Yet, we highlight the striking quantitative distance between the SU(2) and the SU(4) scenarios, with the latter effectively closer to the large-NN limit than to the highly singular SU(2) model.
@@
To further investigate the behavior of correlations in the small-NN regime, we showed how introducing a suitable Raman term in the $N=4$ case induces a crossover between the SU(44) and the SU(22)-symmetric Hubbard model for the Raman-free flavors, while the Raman-coupled ones become completely polarized. This provides a remarkable example of flavor-selective behavior: the Raman-free flavors undergo an SU(22) Mott transition, whereas the Raman-coupled ones form an uncorrelated band insulator. The inter-flavor mutual information neatly captures this crossover, and provides clear-cut insight into the contrasting nature of Raman-coupled and Raman-free flavors. Thus, we reinforce previous evidence that this quantity, as well as closely related geometrical measures of entanglement and quantum correlations, represent a crucial tool not only for the generic detection and characterization of critical behavior and phase transitions, but also for the classification of fermionic states in terms of strong, interaction-induced, correlations [56, 29, 30].
Outlook: Cold-atom experiments and Solid-state counterparts We conclude by briefy discussing how this model, and the tuning from SU(4) to SU(2) are readily realizable in experiments with ultracold atoms in optical lattices, which have repeatedly proven their role as essential platforms to probe the physical properties of SU(NN) fermions [11, 12, 13, 15, 10, 14]. The set-up that we propose can be realized via a simple extension of that of Ref. [10]: starting, for instance, from fermionic 173Yb atoms, one can filter an initial state with equal number of fermions in four nuclear spin components, realizing a SU(4) system which can be loaded in a a three-dimensional optical lattice. The SU(4) symmetry can be broken as in our model by two-photon Raman processes with Rabi frequency $\Omega/h$ (where $h$ is the Planck constant) connecting two flavors and leading to dressed linear combinations. In order to provide experimental evidence, one should measure flavor-selective occupations and double occupations (or correlators) of the form $\langle n_{\alpha}n_{\beta}\rangle$. The double occupancy has been measured in [10] via one-color photoassociation processes that act only when an atom is in a doubly-occupied lattice site. A small magnetic field allows to make the measurement sensitive to the flavors, providing us with flavor-selective quantities and, consequently, with a measurement of the inter-flavor mutual information.
SU(NN) symmetric models of interacting fermions have also been proposed as simplified descriptions of complex layered materials, like moiré transition-metal dichalcogenides [86, 87], where crossovers between SU(NN) models with different NN are also realized using external tunable fields [88, 89], underscoring the multifaceted relevance of our results for state-of-the-art platforms for quantum technology, and opening opportunities to control and assess the correlation properties of quantum materials and their interplay with broken-symmetry phases. A reduction of correlations by increasing the number of components has been also revealed by the behaviour of the spin susceptibility of two-dimensional gases with long-range Coulomb interaction with valley degeneracy [90].
@@
To go beyond our results, it is essential to extend the analysis to account for nonlocal correlations in finite dimensions [29, 91] and multipartite entanglement, as witnessed by quench dynamics [24] or measured by the mutual information between many natural orbitals [74]. While the former two avenues will surely strengthen the connection with experimental observations, the latter has the potential to deepen the understanding of the intricate connection between traditional quantum resources and the concept of fermionic nongaussianity [45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@ Acknowledgements.
Data Availability Statement
All numerical data supporting the findings reported in this manuscript are openly available in [92].